how do you work out 3(x-5) = 12 and what is the answer?
Found 2 solutions by aseem, Alan3354:
Answer by aseem(9)

you querstion........
3x-15 by multiply by 3 you will get 3x-15
x=15-3
x=12

Answer by Alan3354(69443)

how do you work out 3(x-5) = 12 and what is the answer?
3(x-5) = 12
Divide by 3
x-5 = 4
Add 5
x = 9